A company that makes shopping carts for supermarkets and other stores recently purchased some new equipment that reduces the labor content of the jobs needed to produce the shopping carts. Prior to buying the new equipment, the company used 5 workers, who together produced an average of 80 carts per hour. Workers receive $17 per hour, and machine cost was $40 per hour. With the new equipment, it was possible to transfer one of the workers to another department, and equipment cost increased by $10 per hour while output increased by 6 carts per hour.

a.
Compute labor productivity under each system. Use carts per worker per hour as the measure of labor productivity. (Round your answers to 3 decimal places.)

Before carts per worker per hour
After carts per worker per hour


b. Compute the multifactor productivity under each system. Use carts per dollar cost (labor plus equipment) as the measure. (Round your answers to 3 decimal places.)

Before carts/dollar cost
After carts/dollar cost


c. Comment on the changes in productivity according to the two measures. Round your intermediate calculations to 3 decimal places and final answers to 2 decimal places.

Labor productivity
by %
Multifactor productivity
by %

A manager checked production records and found that a worker produced 136 units while working 40 hours. In the previous week, the same worker produced 84 units while working 30 hours.


a. Compute Current period productivity and Previous period productivity. (Round your answers to 2 decimal places.)

Current period productivity Units / hr
Previous period productivity Units / hr


b. Did the worker's productivity increase, decrease, or remain the same? (Round your intermediate calculations and final answer to 2 decimal places. Omit the "%" sign in your response.)

Worker's productivity           

by %
Answer #1

a) Total carts before = 80

Total carts after = 80 + 6 = 86

Workers after = 5- 1 = 4

labor productivity before = 80 / 5 = 16.000 carts per worker per hour

labor productivity after = 86 / 4 = 21.500 carts per worker per hour

b) Total cost earlier system = 17*5 + 40 = $125

Total cost after new machine = 17*4 + 40+10 = $118

Multifactor productivity before = 80 / 125 = 0.640 carts per dollar

Multifactor productivity after = 86 / 118 = 0.729 carts per dollar

c) % Change in labor productivity =(New productivity - Old productivity) / Old Productivity = (21.500 - 16) / 16 = 34.375%

% Change in multifactor productivity = (0.729-0.640) / 0.640 = 13.906%

Next question a) Current period productivity = 136 /40 = 3.40 units/ hr

Previous period productivity = 84 / 30 = 2.80 units/hr

Increase in worker's productivity = (3.40 -2.80) / 2.80 =21.43%